The flywheel of an engine has a moment of inertia 3.00 kg m^2 about its rotation axis.

What constant torque is required to bring it up to an angular speed of 370 rev/min in a time of 7.60 s, starting from rest?


       N?m     

What is its final kinetic energy?

      J
Answer #1

here we have given with values

n = 370 rpm = 370/60 = 6.1667 rev/sec

? = 2?n =2 * 3.14 * 6.1667=38.72

we know that

L = I? = 3 * 38.72 = 116.18 N-m-s

we know that

? = ?L/?t = 116.18/ 7.60=15.2868 N-m

part b ) kinetic energy = Iw^2/2 = 0.5 * 3 * (38.72 ) ^2 = 2248.8576